OpenStudy (anonymous):

Express y in terms of x ln(y)=x+2

OpenStudy (lgbasallote):

divide both sides by x \[\implies \ln y = \frac{x+2}{x}\] change to log form \[\implies \log _e y = \frac{x+2}{x}\] change to exponential form \[\implies y = e^{\frac{x+2}{x}}\] does that help?

OpenStudy (anonymous):

X isn't on the side of the ln(y) though....

OpenStudy (anonymous):

Its simply ln(y) = x+2

OpenStudy (lgbasallote):

oh

OpenStudy (anonymous):

Same things apply here: